OpenStudy (anonymous):

The center pole of a tent is 8 feet long ,and a side of the tent is 12 feet long. If a right triangle is formed where the center pole meets the ground, what is the measure of the missing angle to the nearest degree? a) 34 b) 42 c) 48 d) 56

OpenStudy (anonymous):

cos of the angle = 8/12 angle = 48 c)

OpenStudy (anonymous):

ok with that?

myininaya (myininaya):

how do we know we aren't missing the other angle sinx=8/12 giving us x=42

myininaya (myininaya):

there are two angles right? which one is considered the missing one?

OpenStudy (anonymous):

yes - u r right the missing angle could be either - i dont know why i didnt notice that!!

OpenStudy (anonymous):

i think i need to go to bed

myininaya (myininaya):

lol. don't go

myininaya (myininaya):

i will give you a medal

myininaya (myininaya):

im jk you go i won't hold you back

OpenStudy (anonymous):

well its late here in uk - 12.30 am - thnx for the medal. good night

myininaya (myininaya):

later